# Flattery, Fluff, and Fog: Diagnosing and Mitigating Idiosyncratic Biases in Preference Models

### v2

@@ Abstract
Language models serve as proxies for human preference judgements in alignment and evaluation, yet they exhibit systematic miscalibration, prioritizing superficial patterns over substantive qualities. This bias manifests as overreliance on features like length, structure, and style, leading to issues like reward hacking and unreliable evaluations. Evidence suggests these biases originate in artifacts in human training data. In this work, we systematically investigate the relationship between training data biases and preference model miscalibration across five idiosyncratic features of language model generations: length, structure, jargon, sycophancy and vagueness. Using controlled counterfactual pairs, we first quantify the extent to which preference models favor responses with magnified biases (skew), finding this preference occurs in $>60\%$ of instances, and model preferences show high miscalibration ($\approx 40\%$) compared to human preferences. Notably, bias features only show mild negative correlations to human preference labels (mean $r_{\mathrm{human}}=-0.12$) but show moderately strong positive correlations with labels from a strong reward model (mean $r_{\mathrm{model}}=+0.36$), suggesting that models may overrely on spurious cues.
To mitigate these issues, we propose a simple post-training method based on counterfactual data augmentation (CDA) using synthesized contrastive examples. Fine-tuning models with CDA reduces average miscalibration from $39.4\%$ to $32.5\%$ and average absolute skew difference from $20.5\%$ to $10.0\%$, while maintaining overall RewardBench performance, showing that targeted debiasing is effective for building reliable preference models.111 Our code and data are available at https://www.github.com/anirudhb123/preference-model-biases.
[…]
tcb@breakable
Flattery, Fluff, and Fog:
Diagnosing and Mitigating Idiosyncratic Biases in Preference Models
Anirudh Bharadwaj[Uncaptioned image]  Chaitanya Malaviya[Uncaptioned image]  Nitish Joshi[Uncaptioned image]  Mark Yatskar[Uncaptioned image] [Uncaptioned image]University of Pennsylvania     [Uncaptioned image]New York University {anirudh2,cmalaviy}@seas.upenn.edu
@@ 1 Introduction
Language models are increasingly used as proxies for human preference judgements, both as reward models for aligning models via reinforcement learning from human feedback (RLHF; Stiennon et al., 2020; Ouyang et al., 2022) and as automated evaluators for judging model outputs Zheng et al. (2023). While these preference models serve as a cheap and scalable alternative to human annotation, recent evidence suggests that they can exhibit systematic miscalibration, where they prioritize undesirable or superficial patterns over substantive qualities valued by humans Li et al. (2024).
Prior work has shown that this miscalibration can manifest as overreliance on non-meaningful features such as response length, style, and formatting. For instance, models prefer verbose or list-formatted responses disproportionately Li et al. (2024). Such biases can propagate into downstream applications with undesirable consequences. When used as reward models, they incentivize reward hacking where models optimize for proxy features (e.g., verbosity) that diverge from human preferences Skalse et al. (2022); Chakrabarty et al. (2025). As evaluators, they can distort evaluation conclusions and risk optimizing towards surface-level properties Feuer et al. (2025); Wu and Aji (2025).
These risks are compounded by evidence that biases in preference models may originate from training data artifacts Bansal et al. (2024). Prior work has found correlations between response length and preference labels in preference datasets Singhal et al. (2024), as well as annotators’ stylistic preferences. However, existing studies have primarily documented individual biases in isolation, leaving a gap in quantifying how training data artifacts translate to model miscalibration across various bias dimensions. Crucially, this involves measuring the divergence of model-human preferences when bias features are experimentally isolated.
In this work, we systematically investigate the relationship between training data biases and preference model miscalibration. We focus on five idiosyncratic bias features frequently observed in LM-generated text (§2, also showcased in Figure LABEL:fig:fff): length (verbosity), structure (e.g., list formatting), jargon (overly technical language), sycophancy (excessive user agreement) and vagueness (lack of specificity). To measure model reliance on these features in a controlled manner, we construct counterfactual response pairs where a base response is perturbed to amplify the target bias feature while preserving other meaningful features (e.g., lengthening a concise answer with redundant phrases). Generating these counterfactual pairs for a diverse set of queries, we quantify two metrics: (1) the skew in preference model preferences toward biased responses, and (2) the miscalibration rate: the divergence between model and human preferences on these pairs.
@@
To trace these biases to the training data, we analyze the training data of widely used reward models Liu et al. (2024)4). We find a noticeable imbalance in the presence of biases in human-chosen vs. rejected responses, which can allow reward models to rely on these bias features. Indeed, correlation analysis shows overreliance on bias features: features that are only weakly correlated with human preferences (e.g., $r=-0.09$ for length) are strongly predictive of trained models’ preferences ($r=0.67$). This suggests that standard RLHF pipelines inadvertently magnify subtle data artifacts into misaligned preference signals.
To mitigate these issues, we propose a simple post-training method using counterfactual data augmentation (CDA), which synthesizes contrastive examples to penalize biased preferences (§5). For each bias feature, we augment existing preference datasets with flipped pairs where the perturbed response is explicitly dispreferred. Fine-tuning reward models on this counterfactual data reduces average miscalibration by 6.9% and average absolute skew difference by 10.6%, while maintaining competitive performance on RewardBench Lambert et al. (2025). Our results demonstrate that targeted debiasing of reward models is largely effective with existing alignment pipelines.
@@ 2.1 Problem Formulation
$\uparrow$
Preferred by Human Majority
$\uparrow$
Preferred by Preference Model
Misalignment ($\times$)
@@
    where $s$ represents $R$’s response quality. In RLHF, this reward model can then be used to drive policy updates (e.g., PPO Schulman et al. (2017)) based on the Bradley–Terry model Bradley and Terry (1952). Under this model, the probability of preferring response $R_{1}$ over $R_{2}$ is given by:
@@ 2.2 Biases Under Consideration
    Length: Preference models often favor longer responses, even when the added length doesn’t contribute substantive information Singhal et al. (2024); Dubois et al. (2024). This bias may stem from a training data heuristic, where length is a correlate of comprehensiveness. As a result, models may generate overly verbose responses.
@@
    Structure: Preference models may disproportionately favor responses with bullet lists or numbered points over narrative prose, even when prose is more suitable Li et al. (2024). This could be learned if structured formats are overrepresented in responses preferred by annotators. The consequence is a potential overuse of listicles, leading to outputs that feel formulaic or fail to convey arguments that benefit from prose.
@@
    Jargon: This refers to a preference for responses using specialized or domain-specific terminology even when it is not necessary. Models might learn this if the presence of jargon in the training data is correlated with highly-preferred responses, leading them to use it as a proxy for quality. Resultingly, models may generate responses that give a superficial impression of expertise without being more useful.
@@
    Sycophancy: This involves the model agreeing with or validating the user’s stated opinions and assumptions, rather than offering a neutral and objective response Sharma et al. (2024); Perez et al. (2023). This behavior may stem from training data if sycophantic responses were more often preferred by human annotators. The downside of this bias is that models may reinforce a user’s biases, fail to provide objective information and appear less trustworthy.
@@
    Vagueness: This bias is characterized by models favoring responses that make broad statements that cover multiple aspects superficially, rather than providing concrete information that specifically addresses the query (example in Table 2). This may stem from vague statements being less falsifiable, and thus less penalized in training data. Such vague outputs can lead to responses that are unhelpful and lack depth.
@@ 2.3 Counterfactual Testing
To approximate this perturbation function $f_{p}$, we use the RATE (Rewrite-based Attribute Treatment Estimators) protocol Reber et al. (2025). Specifically, we first use a language model to rewrite an original base response to produce a counterfactual response $R^{\prime}_{p}$ that amplifies the bias feature $p$. We then correct for the rewriting error by rewriting again to produce a new base response $R_{p}$. Using the pair $(R_{p},R^{\prime}_{p})$, we compute $W_{RM}(Q,R_{p})$ and $W_{RM}(Q,R^{\prime}_{p})$ to measure the causal effect of $p$ on the reward score. Similarly, we compute $W_{\textnormal{EVAL}}(Q,R_{p},R^{\prime}_{p})$ to measure the effect of $p$ on $W_{\textnormal{EVAL}}$’s evaluation judgment.
@@
    Miscalibration Rate: This measures the degree of disagreement between the model’s preference and the aggregated human majority preference for $R^{\prime(i)}_{p}$ over $R^{(i)}_{p}$ (denoted as $\text{Human}(R^{\prime(i)}_{p}>R^{(i)}_{p})$).
@@ 3 Are preference models misaligned with human judgements?
Figure 2: Contingency tables for each bias feature in the $2500$-example training subset, showing co-occurrence of bias presence in human-chosen vs. human-rejected responses. Anti-diagonal cells (top-right and bottom-left) quantify cases where the two responses differed on the feature.
@@
To study structure, jargon, and length biases, we sample 100 queries from Chatbot Arena Chiang et al. (2024), where we filter single-sentence queries that are in English, grammatically well-formed (ending with a question mark), semantically meaningful and non-offensive. Since sycophancy is plausible when the user expresses an opinion, we generate 100 queries that express an opinion (e.g., "Isn’t modern art just lazy compared to classical techniques?"). To study vagueness, we use 78 human-written NLP-related queries from the KIWI dataset Xu et al. (2024), and generate 22 additional queries that cover similar content. This choice is motivated by the fact that scientific queries inherently require unambiguous and detailed responses and a preference for vague responses would represent a clear flaw. For all these queries, we generate counterfactual pairs of responses using the RATE protocol. Prompts are provided in Appendix B.
@@
We consider four reward models that are all trained on 80k preference pairs from the Skywork reward data collection Liu et al. (2024): Gemma2-2B, Gemma-2-27B, Llama-3.1-8B, Llama3.2-3B. In addition, we consider three proprietary models that we use as LLM evaluators: Gemini-2.5-Pro Team et al. (2023), GPT-4o Hurst et al. (2024) and Claude-3.7-Sonnet Anthropic (2025). For response generation, we used GPT-4o.
@@ Results.
Figure 1 shows that LLM evaluators also exhibit significant miscalibration relative to human preferences, with the largest deviations in length and vagueness biases. LLM evaluators similarly amplify skew toward perturbed responses compared to human annotators, particularly for vagueness and sycophancy. Notably, LLM evaluators show a dramatically higher preference for sycophantic responses ($\sim$75-85skew) compared to humans ($\sim$50%). These findings reveal that preferences of LLM evaluators can similarly diverge from human preferences.
@@ 4 Are preference models over-reliant on biases in training data?
Each bias is first mapped to a measurable quantity that can be automatically extracted from a response; for example, for length we record the token-count of the response, whereas for structure we use a binary flag that detects the presence (or absence) of list-style formatting. The exact labeling heuristics used for all biases are provided in Appendix C. For each query-response pair $(Q,R_{1},R_{2})$ in the perturbation data corresponding to each bias, we then compute feature values for both responses $f_{1}$ and $f_{2}$. We begin by examining how often bias features appear in human-preferred responses within a $2500$-example subset of the training data. Figure 2 presents contingency tables that count these co-occurrences for each bias feature in the human-chosen versus human-rejected responses. The anti-diagonal cells are particularly revealing, as they quantify cases where the two responses differed on that feature.
@@
To quantify the relationship between these biases and preferences more formally, we conduct a correlation analysis. We use both our perturbed counterfactual data and the $2500$-example training data subset. Each bias is similarly mapped to a measurable quantity and for each query-response pair $(Q,R_{1},R_{2})$, we compute feature values $f_{1},f_{2}$ for the responses and their difference $\Delta f$. We then calculate three point-biserial correlations Lev (1949):
@@ 5 Counterfactual Data Augmentation
Figure 5: Average RewardBench scores before and after fine-tuning on counterfactual examples, by bias type. Error bars show ±$\pm 1$ standard error.
@@
We start with the original Skywork training corpus and label both the chosen ($R_{\textnormal{chosen}}$) and rejected ($R_{\textnormal{rejected}}$) responses for the presence of the target bias $p$ (using prompts in Appendix C for some biases and automated methods for others like length). For pairs where neither $R_{\textnormal{chosen}}$ or $R_{\textnormal{rejected}}$ exhibits the bias, we synthesize a new, biased version of the rejected response, $R_{\textnormal{rejected,p}}$. This is done by prompting GPT-4o with "rewrite" instructions (see Prompts B-B) designed to amplify only the feature $p$ in $R_{\textnormal{rejected}}$. This results in new counterfactual instances of the form $(Q,R_{\textnormal{chosen}}\succ R_{\textnormal{rejected,p}})$. In these new pairs, the original chosen response is explicitly preferred over a version of the rejected response that now exhibits the undesired bias. To mitigate potential distribution shifts from these examples, we supplement the counterfactual data with additional examples sampled proportionally from Chatbot Arena (see Table  4). Finally, base reward models are finetuned on this augmented dataset.

### v3

@@ Abstract
Language models serve as proxies for human preference judgements in alignment and evaluation, yet they exhibit systematic miscalibration, prioritizing superficial patterns over substantive qualities. This bias manifests as overreliance on features like length, structure, and style, leading to issues like reward hacking and unreliable evaluations. However, the connection between training data artifacts and the miscalibrated preferences exhibited by models remains poorly understood.
In this work, we systematically investigate the relationship between training data biases and preference model miscalibration across five idiosyncratic features of language model generations: length, structure, jargon, sycophancy and vagueness. Using controlled counterfactual pairs, we first quantify the extent to which preference models favor responses with artificially magnified biases (skew), finding this preference occurs in $60\%$ of instances, and model preferences show high miscalibration ($\approx 40\%$) compared to human preferences. Notably, bias features only show mild negative correlations to human preference labels (mean $r_{\mathrm{human}}=-0.12$) but show moderately strong positive correlations with labels from a strong reward model (mean $r_{\mathrm{model}}=+0.36$), suggesting that models may overrely on spurious cues.
To mitigate these issues, we propose a simple post‐training method based on counterfactual data augmentation (CDA) using synthesized contrastive examples. Fine‐tuning models with CDA reduces average miscalibration from $39.4\%$ to $32.5\%$ and average absolute skew difference from $20.5\%$ to $10.0\%$, while maintaining overall RewardBench performance, indicating that targeted debiasing can strengthen the reliability of preference models within standard alignment pipelines.  [^1] Our code and data are available at https://github.com/anirudhb123/preference-model-biases.
@@ 1 Introduction
Language models are increasingly used as proxies for human preference judgements, both as reward models for aligning models via reinforcement learning from human feedback (RLHF; Stiennon et al., 2020; Ouyang et al., 2022) and as automated evaluators for judging model outputs (Zheng et al., 2023). While these preference models serve as a cheap and scalable alternative to human annotation, recent evidence suggests that they can exhibit systematic miscalibration, where they prioritize undesirable or superficial patterns over substantive qualities valued by humans (Li et al., 2024).
Prior work has shown that this miscalibration can manifest as overreliance on non-meaningful features such as response length, style, and formatting. For instance, models prefer verbose or list-formatted responses disproportionately (Li et al., 2024). Such biases can propagate into downstream applications with undesirable consequences. When used as reward models, they incentivize reward hacking where models optimize for proxy features (e.g., verbosity) that diverge from human preferences (Skalse et al., 2022; Chakrabarty et al., 2025). As evaluators, they can distort evaluation conclusions and risk optimizing towards surface-level properties (Feuer et al., 2025; Wu & Aji, 2025).
These risks are compounded by evidence that biases in preference models may originate from training data artifacts (Bansal et al., 2024). Prior work has found correlations between response length and preference labels in preference datasets (Singhal et al., 2024), as well as annotators’ stylistic preferences. However, existing studies have primarily documented individual biases in isolation, leaving a gap in quantifying how training data artifacts translate to model miscalibration across various bias dimensions. Crucially, this involves measuring the divergence of model-human preferences when bias features are experimentally isolated.
Figure 1: Examples of three idiosyncratic biases in language models: (1) Flattery: responses that excessively agree with the user; (2) Fluff: verbose, uninformative responses; and (3) Fog: vague responses that state many non-specific claims. Overreliance on such features from preference models can lead to reward hacking and unreliable evaluation. The complete list of biases explored in this work is in Table 1.
@@
In this work, we systematically investigate the relationship between training data biases and preference model miscalibration. We focus on five idiosyncratic bias features frequently observed in LM-generated text (§2, also showcased in Figure 1): length (verbosity), structure (e.g., list formatting), jargon (overly technical language), sycophancy (excessive user agreement) and vagueness (lack of specificity). To measure model reliance on these features in a controlled manner, we construct counterfactual response pairs where a base response is perturbed to amplify the target bias feature while preserving other meaningful features (e.g., lengthening a concise answer with redundant phrases). Generating these counterfactual pairs for a diverse set of queries, we quantify two metrics: (1) the skew in preference model preferences toward biased responses, and (2) the miscalibration rate: the divergence between model and human preferences on these pairs.
@@
To trace these biases to the training data, we analyze the training data of widely used reward models (Liu et al., 2024)4). We find a noticeable imbalance in the presence of biases in human-chosen vs. rejected responses, which can allow reward models to rely on these bias features. Indeed, correlation analysis shows overreliance on bias features: on average, they are nearly three times more predictive of trained models’ preferences ($r=+0.36$) than of human preferences (r=0.12r=-0.12). This suggests that standard RLHF pipelines inadvertently magnify subtle data artifacts into misaligned preference signals.
To mitigate these issues, we propose a simple post‐training method using counterfactual data augmentation (CDA), which synthesizes contrastive examples to penalize biased preferences (§5). For each bias feature, we augment existing preference datasets with flipped pairs where the perturbed response is explicitly dispreferred. Fine‐tuning reward models on this counterfactual data reduces average miscalibration by 6.9% and average absolute skew difference by 10.6%, while maintaining competitive performance on RewardBench (Lambert et al., 2025). Our results demonstrate that targeted debiasing of reward models is largely effective with existing alignment pipelines.
@@ 2.1 Problem Formulation
    where $s$ represents $R$’s response quality. In RLHF, this reward model can then be used to drive policy updates (e.g., PPO (Schulman et al., 2017)) based on the Bradley–Terry model (Bradley & Terry, 1952). Under this model, the probability of preferring response $R_{1}$ over $R_{2}$ is given by:
@@ 2.2 Biases Under Consideration
    Length: Preference models often favor longer responses, even when the added length doesn’t contribute substantive information (Singhal et al., 2024; Dubois et al., 2024; Shen et al., 2023). This bias may stem from a training data heuristic, where length is a correlate of comprehensiveness. As a result, models may generate overly verbose responses.
@@
    Structure: Preference models may disproportionately favor responses with bullet lists or numbered points over narrative prose, even when prose is more suitable (Li et al., 2024). This bias may stem from structured formats being overrepresented in responses preferred by annotators. The consequence is a potential overuse of listicles, leading to outputs that feel formulaic or fail to convey arguments that benefit from prose. Prior work has further shown that such format biases, including preferences for lists, emojis, and other stylistic markers, are pervasive in preference models and can be easily amplified during alignment (Zhang et al., 2025).
@@
    Jargon: This refers to a preference for responses using specialized or domain-specific terminology, even when it is not necessary. Models might learn this if the presence of jargon in the training data is correlated with highly preferred responses, leading them to use it as a proxy for quality. Resultingly, models may generate responses that give a superficial impression of expertise without being more useful.
@@
    Sycophancy: This involves the model agreeing with or validating the user’s stated opinions and assumptions, rather than offering a neutral and objective response (Sharma et al., 2024; Perez et al., 2023). This behavior may arise from training data if human annotators preferred sycophantic responses more often. The downside of this bias is that models may reinforce a user’s biases, fail to provide objective information, and appear less trustworthy.
@@
    Vagueness: This bias is characterized by models favoring responses that make broad statements that cover multiple aspects superficially, rather than providing concrete information that specifically addresses the query (example in Appendix Table 2). This may stem from vague statements being less falsifiable, and thus less penalized in training data. Such vague outputs can lead to responses that are unhelpful and lack depth.
@@ 2.3 Counterfactual Testing
To approximate this perturbation function $f_{p}$, we use the RATE (Rewrite-based Attribute Treatment Estimators) protocol (Reber et al., 2025). Specifically, we first use a language model to rewrite an original base response to produce a counterfactual response $R^{\prime}_{p}$ that amplifies the bias feature $p$. We then correct for the rewriting error by rewriting again to produce a new base response $R_{p}$. Using the pair $(R_{p},R^{\prime}_{p})$, we compute $W_{RM}(Q,R_{p})$ and $W_{RM}(Q,R^{\prime}_{p})$ to measure the causal effect of $p$ on the reward score. Similarly, we compute $W_{\textnormal{EVAL}}(Q,R_{p},R^{\prime}_{p})$ to measure the effect of $p$ on $W_{\textnormal{EVAL}}$’s evaluation judgment.
@@
[…]
@@ 3 Are preference models misaligned with human judgements?
To study structure, jargon, and length biases, we sample 100 queries from Chatbot Arena (Chiang et al., 2024), where we filter single-sentence queries that are in English, grammatically well-formed (ending with a question mark), semantically meaningful and non-offensive. Since sycophancy is plausible when the user expresses an opinion, we generate 100 queries that express an opinion (e.g., "Isn’t modern art just lazy compared to classical techniques?"). To study vagueness, we use 78 human-written NLP-related queries from the KIWI dataset (Xu et al., 2024), and generate 22 additional queries that cover similar content. Scientific queries inherently require unambiguous and detailed responses and a preference for vague responses would represent a clear flaw. For all these queries, we generate counterfactual pairs of responses using the RATE protocol. Prompts are provided in Appendix B.
@@
We consider four reward models, all trained on v0.2 of the Skywork reward data collection (Liu et al., 2024), which aggregates diverse preference datasets (e.g., HelpSteer2, OffsetBias, WildGuard) and underlies many top-performing open-source reward models. Specifically, we study Gemma2-2B, Gemma-2-27B, Llama-3.1-8B, and Llama-3.2-3B. In addition, we consider three proprietary models that we use as LLM evaluators: Gemini-2.5-Pro (Team et al., 2023), GPT-4o (Hurst et al., 2024), and Claude-3.7-Sonnet (Anthropic, 2025). For response generation, we used GPT-4o.
@@ Results.
Figure 2 shows that LLM evaluators also exhibit significant miscalibration relative to human preferences, with the largest deviations in length and vagueness biases. LLM evaluators similarly amplify skew toward perturbed responses compared to human annotators, particularly for vagueness and sycophancy. Notably, LLM evaluators show a dramatically higher preference for sycophantic responses ($\sim$75-85% skew) compared to humans ($\sim$50%). These findings reveal that preferences of LLM evaluators can similarly diverge from human preferences.
@@ 4 Are preference models over-reliant on biases in training data?
Figure 3: Contingency tables for each bias feature in the 25002500‑example training subset, showing co‑occurrence of bias presence in human‑chosen vs. human‑rejected responses. Anti-diagonal cells (top-right and bottom-left) quantify cases where the two responses differed on the feature.
@@
Each bias is first mapped to a measurable quantity that can be automatically extracted from a response; for example, for length we record the token-count of the response, whereas for structure we use a binary flag that detects the presence (or absence) of list‑style formatting. The exact labeling heuristics used for all biases are provided in Appendix C. For each query-response pair $(Q,R_{1},R_{2})$ in the perturbation data corresponding to each bias, we then compute feature values for both responses $f_{1}$ and $f_{2}$. We begin by examining how often bias features appear in human-preferred responses within a $2500$-example subset of the training data. Figure 3 presents contingency tables that count these co-occurrences for each bias feature in the human-chosen versus human-rejected responses. The anti-diagonal cells are particularly revealing, as they quantify cases where the two responses differed on that feature.
@@
To quantify the relationship between these biases and preferences more formally, we conduct a correlation analysis. We use both our perturbed counterfactual data and the $2500$-example training data subset. Each bias is similarly mapped to a measurable quantity and for each query-response pair $(Q,R_{1},R_{2})$, we compute feature values $f_{1},f_{2}$ for the responses and their difference $\Delta f$. We then calculate three point-biserial correlations (Lev, 1949):
@@ 5 Counterfactual Data Augmentation
We start with the original Skywork training corpus and label both the chosen ($R_{\textnormal{chosen}}$) and rejected ($R_{\textnormal{rejected}}$) responses for the presence of the target bias $p$ (using prompts in Appendix C for some biases and automated methods for others like length). For pairs where neither $R_{\textnormal{chosen}}$ or $R_{\textnormal{rejected}}$ exhibits the bias, we synthesize a new, biased version of the rejected response, $R_{\textnormal{rejected,p}}$. This is done by prompting GPT-4o with "rewrite" instructions (see Prompts B.1-B.1) designed to amplify only the feature $p$ in $R_{\textnormal{rejected}}$. This results in new counterfactual instances of the form $(Q,R_{\textnormal{chosen}}\succ R_{\textnormal{rejected,p}})$. In these new pairs, the original chosen response is explicitly preferred over a version of the rejected response that now exhibits the undesired bias. To mitigate potential distribution shifts from these examples, we supplement the counterfactual data with additional examples sampled proportionally from Chatbot Arena (see Table  4). Finally, base reward models are finetuned on this augmented dataset.